OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation k and the inverse variation equation for the situation. y=4 when x=5 Find the constant of variation k. k= Write the inverse variation equation. =

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (anonymous):

Allsia1245 what?

OpenStudy (sleepyhead314):

y = k / x plug in y = 4 and x = 5 and solve for k

OpenStudy (anonymous):

4=k/5

OpenStudy (sleepyhead314):

solve for k now :) multiply both sides by 5

OpenStudy (anonymous):

Sorry @Novice3000 it was my friend that typed this

OpenStudy (anonymous):

k=20?

OpenStudy (sleepyhead314):

yes

OpenStudy (anonymous):

it's alright lol not mad lol

OpenStudy (sleepyhead314):

then y = k / x plug in the "k" and keep the y and x to get the equation

OpenStudy (anonymous):

ah so 4=20/5 is the equation for the last one

OpenStudy (sleepyhead314):

nope keep the y and x variables in there the 4 and 5 were just temporary :)

OpenStudy (anonymous):

oh so y=20/x

OpenStudy (sleepyhead314):

yes :)

OpenStudy (anonymous):

Yes, thank you again :D